Estimate the number of impacts that air molecules make with the palm of your hand each second, assuming that all air consists entirely of nitrogen (molar mass=0.028kg)?



1
Expert's answer
2021-11-03T09:50:35-0400

Force= pressure"\times" area

Force = number of hits mass change "\times" velocity Change

Number of hits ="\frac{air pressure\times A}{m}"

Number of hits ="\frac{1 atm \times A}{0.028}"